OpenStudy (sissyedgar):

Help please, Medals! :) Serena bought 3/4 pound of cheese. she used 2/3 or cheese for sandwitches. How much cheese did Serena use for sandwiches? A. 1/4lb B. 1/3lb C. 1/2lb D. 5/7lb

OpenStudy (ultrasoundwave15):

Try this :) http://openstudy.com/study#/updates/529516bfe4b0999412bdd844

OpenStudy (ultrasoundwave15):

I believe it's 1/2

OpenStudy (sissyedgar):

mkay, well ty!

OpenStudy (bookworm14):

It would be \[\frac{ 3 }{ 4 } * \frac{ 2 }{ 3 }=\frac{ 1 }{ 2 }\] you just multiply across 3*2=6 4*3=12 6/12= 1/2 Hope this helps~Bree :)

OpenStudy (sissyedgar):

Ty Bree!

OpenStudy (sissyedgar):

U got 100%

OpenStudy (sissyedgar):

I*